OpenStudy (anonymous):

help me plz Choose the slope-intercept equation of the line that passes through the point (4, -6) and is parallel to y = - x + 5

OpenStudy (cwrw238):

standard form of slope intercept line is y = mx + b where m = slope and b = y-intercept

OpenStudy (cwrw238):

now as its parallel to y = - x + 5 the lines slope is the same so what is the slope of y = -x + 5 - compare it with the standard equation y = mx + b

OpenStudy (anonymous):

A.y = -2x + 2 B.y = 2x - 14 C.y = x - 8 D.y = - x - 4

OpenStudy (cwrw238):

m is the slope

OpenStudy (anonymous):

is it D

OpenStudy (cwrw238):

i dont know - but im trying to guide you through solving the problem